OpenStudy (jojo12g):

Add. 51.3 + (–51.3) A. –102.6 B. –51.3 C. 0 D. 102.6

OpenStudy (alexandervonhumboldt2):

51.3+(-51.3)=51.3-51.3=??? +(-)=-

OpenStudy (ddos_dragon):

0

OpenStudy (ddos_dragon):

Why?

OpenStudy (alexandervonhumboldt2):

a-a=0

OpenStudy (ddos_dragon):

If a negative answer comes across its parallel positive number.

OpenStudy (ddos_dragon):

It will end up equaling 0.

OpenStudy (jojo12g):

thanks guys

OpenStudy (ddos_dragon):

No problem.